### v2

@@ 1 Introduction
In [1] the authors studied this process on several graphs, and established results on the dispersion time and the maximum distance of any particle from the origin at the point in time of dispersion. A particular case considered in [1] is the behaviour when the underlying graph is the complete graph $K_{n}$ with $n$ vertices. In that case we denote the dispersion time by $T_{n,M}$, since, due to symmetry, the choice of the initial vertex does not matter. The most general results come from considering a lazy variant of the dispersion process, which was shown to disperse more particles in a comparable number of steps. More precisely, in this lazy version any unhappy particle moves with probability $q\in(0,1]$ and stays at its current location with probability $1-q$. The main result of [1] regarding $T_{n,M}$ is that there are constants $c,C>0$ such that, if $M=(1-q/2-\alpha)n$ where $\alpha>0$ is allowed to depend on $n$, then
$T_{n,M}\leq C(q\alpha)^{-1}\ln n~{}\text{ with probability at least }~{}1-C/n,$ (1)
@@
$T_{n,M}\geq e^{cnq^{3}\alpha^{4}}~{}\text{ with probability at least }~{}1-e^{-cnq^{3}\alpha^{4}}.$ (2)
@@
Our main contribution is a thorough analysis of the dispersion process when, as above, we assume that $M=(1+\varepsilon)n/2$ and $|\varepsilon|=o(1)$. We establish that the process exhibits three qualitatively different behaviours based on the asymptotics of $\varepsilon$, where, informally speaking, $T_{n,M}$ smoothly changes from $|\varepsilon|^{-1}\ln(\varepsilon^{2}n)$ to $n^{1/2}$ and then to $\varepsilon^{-1}e^{\Theta(\varepsilon^{2}n)}$; in particular, $T_{n,M}=\Theta(n^{1/2})$ whenever $M=n/2+O(\sqrt{n})$ within a scaling window of size $O(\sqrt{n})$. We continue with a detailed and formal description of the dispersion process and state our results.
@@
These results imply that there is critical window around $n/2$ of order $n^{1/2}$, where the dispersion time increases smoothly from logarithmic to $n^{1/2}$ and then to exponential. Apart from these estimates we can also use our main theorems to obtain information about, for example, the expectation of $T_{n,M}$. In particular, Theorem 1.1 guarantees that $T_{n,M}$ has an exponential(-ly thin) upper tail and so $T_{n,M}$ is, after appropriate normalization, integrable; we readily obtain that
@@ 1.1 Main Proof Ideas
So, as long as $U_{t}$ is (much) larger than $|\varepsilon|n$, then $U_{t+1}$ is (much) smaller than $U_{t}$ in expectation. In other words, when there are ‘many’ unhappy particles at time $t$, the expected number of $U_{t+1}$ decreases by a considerable amount. However, this is no longer true when $U_{t}=O(|\varepsilon|n)$: then the expected decrease of $U_{t}$ might be tiny, or when $\varepsilon>0$, we can even expect that the number of unhappy particles increases. We will use different methods to study the trajectory of $U_{t}$, depending on the range of $\varepsilon$ and whether we are considering the upper or the lower tail.
@@
Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Given $k\in\mathbb{N}$, we write $[k]\coloneqq\{1,\dots,k\}$. Given functions $f:\mathbb{N}\mapsto\mathbb{R},g:\mathbb{N}\mapsto\mathbb{R}$, we write either $f\ll g$ or $f=o(g)$ if $f(n)/g(n)\rightarrow 0$ as $n\rightarrow\infty$ and $f=O(g)$ if there is a constant $C>0$ such that $f(n)\leq Cg(n)$ for all large enough $n$. We write $f=\Theta(g)$ if $f=O(g)$ and $f=\Omega(g)$, where the latter is the same as $g=O(f)$. Given random variables $X$ and $Y$, we write $X=_{d}Y$ if $X$ and $Y$ have the same distribution, whereas we write $X\leq_{sd}Y$ if $Y$ stochastically dominates $X$, meaning that $\mathbb{P}(X\geq x)\leq\mathbb{P}(Y\geq x)$ for all $x\in\mathbb{R}$.
@@ 2.2 Bounds on $\mathbb{E}[U_{t+1}\mid U_{t}]$
where the last step can be verified numerically. Next turn to the main case $H_{t}\geq 1$. Using $1+x\leq e^{x}$ for $x\in\mathbb{R}$ and $e^{-x}\leq 1-x+x^{2}/2$ for $x\geq 0$ (in this order) we infer
@@
Moreover, since $H_{t}-1=M-U_{t}-1$ and $M=(1+\varepsilon)n/2$ and $|\varepsilon|<1/9$, we immediately obtain
@@
$\mathbb{E}\big{[}U_{t+1}\mid U_{t}\big{]}\leq U_{t}(1+\varepsilon)-\frac{U^{2}_{t}}{n}.$
@@ 2.3 Reducing the Number of Unhappy Particles
Here we state and prove a lemma that provides information on the number of steps required to reduce the number of unhappy particles from some number $\ell\leq M$ to some $h<\ell$.
@@
Let $n\in\mathbb{N}$, $t_{0}\in\mathbb{N}_{0}$, $|\varepsilon|<1$, $M=(1+\varepsilon)n/2\in\mathbb{N}$ and $h\geq\max\{2\varepsilon n,2\}>0$. Set $\tau_{h}\coloneqq\inf\{t\geq t_{0}:U_{t}<h\}$. Then, for any $b>0$ and for every $\ell\geq h$
@@ 2.4 Branching Process Approximation
In this subsection we introduce a process that we will use in order to obtain upper and lower couplings for the process of unhappy particles. For $|\chi|<1$ and some random variable $N\in\mathbb{N}$ define a stochastic process $(Z_{t}(N,\chi))_{t\in\mathbb{N}_{0}}$ by
$Z_{0}(N,\chi)=N\quad\text{and}\quad Z_{t}(N,\chi)=_{d}2\,\text{Bin}\big{(}Z_{t-1}(N,\chi),(1+\chi)/2\big{)},~{}t\in\mathbb{N}.$ (8)
We call $(Z_{t}(N,\chi))_{t\in\mathbb{N}_{0}}$ the binomial process in this paper. We want to stress that the process is well defined even if $\chi<0$ (as long as $\chi>-1$). When considering $Z_{t}(N,\chi)$ we sometimes drop the dependence on $N$ or on $\chi$, if it is clear from the context or otherwise convenient. In the next lemma we show how to control the number of unhappy particles at some given step in terms of the binomial process.
@@
To see this, at step $t_{0}+t$, remove the particles in ${\cal U}_{t_{0}+t-1}$ and put them back one after the other (by choosing for each one of them independently and uniformly at random a vertex) to obtain ${\cal U}_{t_{0}+t}$. Note that, during this process, the probability that the $j$-th particle, where $1\leq j\leq U_{t_{0}+t-1}$, is put on one of the, at most $M$, currently occupied vertices is at most $M/n$. Moreover, if this event occurs, then the number of unhappy particles increases by at most 2; it increases by exactly 2 if the $j$-th particle is put on a vertex that is currently occupied by exactly one other particle. This establishes (9). With this at hand we proceed by showing the statement of the lemma by induction over $t$. Note that, for $t=0$, we have that $U_{t_{0}}=Z_{0}$. Next, suppose that the statement is true for $t-1$, where $t\in\mathbb{N}$. Then, for any $z\in\mathbb{N}_{0}$,
@@
Since $U_{t_{0}+t-1}\leq\delta n$ by definition of $\tau$, the last expression is at least $(1+\varepsilon)/2-2\delta\eqqcolon q$. Therefore, by induction on $t$, if $(B_{i,t})_{i,t\geq 1}$ denotes a doubly-infinite sequence of iid $\text{Ber}(q)$ random variables,
@@
In order to study the binomial process $(Z_{t}(N,\chi))_{t\in\mathbb{N}_{0}}$ from (8) we provide an alternative description of it. We can think of any particle $p$ counted in $Z_{0}$ to be the root of an independent Galton-Watson tree $\mathcal{T}_{p}$, where every individual has $2\text{Ber}((1+\chi)/2)$ offspring, that is, no offspring with probability $(1-\chi)/2$ and 2 offspring with probability $(1+\chi)/2$. Obviously, $Z_{t}$ equals in distribution the total number of individuals in all $\mathcal{T}_{p}$’s at distance $t$ from the root. Let $\mathcal{T}(\chi)$ be a Galton-Watson tree with the aforementioned distribution, and let $\{\mathcal{T}(\chi)\leq t\}$ be the event that this branching process has at most $t$ levels (in particular, there are no individuals at distance $t$ from the root). Then we immediately obtain the following observation, tailored to our intended application; note that $\varepsilon\in(-1,1)$, i.e., the parameter of the process can be negative, but $(1+\varepsilon)/2>0$ always.
@@
we then obtain (10). The lower bound is established in a similar way. Indeed, we obtain from $y_{k}=\frac{y_{k-1}}{1+\varepsilon}(1-x_{k-1}/2)^{-1}$ (see the first ’$=$’ in (12)) using $\sum_{j=2}^{\infty}q^{j}\leq 2q^{2}$ for all $q\in[0,1/2]$ that
@@ 3 The Upper Tail
Throughout this section we assume that $M=(1+\varepsilon)n/2$, where $|\varepsilon|=o(1)$ and $|\varepsilon|<1/9$ and that, say, $n\geq 100$, so that the statements from the previous section apply. Define
@@
that we will sometimes use without further reference. The main goal of this section is to prove the following statement, from which the bounds stated in Theorem 1.1 follow immediately. No attempt was made to optimize the constants.
@@
Set $I_{i}=[(i-1)k,ik)$ for all $i\in\mathbb{N}$ so that $\mathbb{N}_{0}$ is the union of all those intervals. Note that if $T_{n,M}$ is greater than $Aq^{-1}k$, then necessarily the process of unhappy particles cannot terminate in any of the intervals $I_{i}$ with $i\leq Aq^{-1}$. Denote by $\mathcal{S}_{i}$ the event that the process of unhappy particles does not terminate until the end of $I_{i}$; that is,
@@
Denoting by $J$ the (first, random) index $i\in\mathbb{N}$ such that $\mathcal{S}_{i}$ does not occur, we can bound
$\mathbb{P}(T_{n,M}>Aq^{-1}k)\leq\mathbb{P}(J>\lfloor Aq^{-1}\rfloor)\leq\prod_{1\leq i\leq\lfloor Aq^{-1}\rfloor}\mathbb{P}\big{(}\mathcal{S}_{i}\mid\cap_{r=1}^{i-1}\mathcal{S}_{r}\big{)}.$ (15)
@@
If $k\geq\hat{\varepsilon}^{-1}$ and $q\in\mathbb{R}$ satisfy
@@
Let $\tau\geq t_{0}$ be such that $U_{\tau}<16\hat{\varepsilon}n$. By conditioning on $\tau$ and $U_{\tau}$ and using the Markov property of the process once more we obtain from the two observations that
@@
Recall that $U_{t}=0$ is equivalent to $T_{n,M}\leq t$. Also, recall the definition of the binomial process that we introduced in Section 2.4. Throughout this proof we let $k\coloneqq\lfloor\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)\rfloor$. We obtain, by applying Lemma 2.8 for any $1\leq r<16\hat{\varepsilon}n$,
@@ 4 The Lower Tail
Throughout this section we assume that $M=(1+\varepsilon)n/2$, where $|\varepsilon|=o(1)$ and $|\varepsilon|<1/9$ and that, say, $n\geq 100$, so that the statements from the Section 2 apply. Define
$\hat{\varepsilon}\coloneqq\max\{|\varepsilon|,en^{-1/2}\}$
like in the previous section. We show the following statement, from which Theorem 1.2 follows immediately.
@@
$\mathbb{P}\big{(}T_{n,M}\leq\hat{\varepsilon}^{-1}k_{0}/(2A)\big{)}\leq\left\{\begin{array}[]{ll}3\exp\left(-\frac{A\hat{\varepsilon}^{2}n}{2^{18}k_{0}}\right)&\mbox{, if }A>k_{0}\\ 3A^{-1}&\mbox{, if }1\leq A\leq k_{0}\end{array}\right.\enspace.$
@@
The next lemma asserts that $\mathcal{A}(b)$ occurs for $b=\exp(O(\hat{\varepsilon}n))$ with (very) high probability, meaning that it is very unlikely for $U_{t+1}$ to drop below $2^{-5}U_{t}$ if $U_{t}>2^{-5}\hat{\varepsilon}n$, where $t\leq b$.
@@
The second issue arising from using our coupling from Lemma 2.9 between $U_{t}$ and the binomial process is that it only works over multiple time steps when the number of unhappy particles remains small throughout. In particular, the coupling provides good estimates only when the number of unhappy particles remains below, say, $\delta n$, where $\delta=\hat{\varepsilon}/8$. In order to ensure that there are sufficiently many time steps until the number of unhappy particles goes above $\delta n$, we introduce a second barrier at $\hat{\varepsilon}n/32$, where we initiate the coupling. More specifically, we define $\tau^{U}_{0}\coloneqq 0$ and iteratively, for $i\in\mathbb{N}$,
@@
where the $D_{\ell}$’s are iid such that $D_{\ell}=1$ with probability $(1+\varepsilon)/2$ and $D_{\ell}=-1$ with probability $(1-\varepsilon)/2$. Since the probability that an unhappy particle remains unhappy after moving to a randomly chosen vertex is at most $M/n=(1+\varepsilon)/2$, we note that $U_{\tau^{L}_{i}+1}\leq_{sd}R_{U_{\tau^{L}_{i}}}$ and in general
$U_{\tau^{L}_{i}+k}\leq_{sd}R_{U_{\tau^{L}_{i}}+\cdots+U_{\tau^{L}_{i}+k-1}},\text{ }k\in\mathbb{N}.$
@@
throughout. We assume that $k\geq 1$ and $A\geq 4$, as otherwise there is nothing to show. Note that if $T_{n,M}\leq\tau_{1}^{L}+k$, then either $k\leq\tau_{1}^{U}-\tau_{1}^{L}$ and the process stops within $k$ steps, or $k>\tau_{1}^{U}-\tau_{1}^{L}$. Since, by Lemma 2.9, the binomial process $Z_{t}(U_{\tau_{1}^{L}},\varepsilon-\hat{\varepsilon}/2)$ provides a lower coupling for $U_{\tau_{1}^{L}+t}$ until time $\tau_{1}^{U}$, we obtain
@@
In order to proceed note that $x\mapsto x/\ln x$ is increasing for $x\geq e^{2}$ and thus
$\hat{\varepsilon}^{2}n/\ln(\hat{\varepsilon}^{2}n)\geq e^{2}/2.$ (26)
@@
This means that $\mathbb{P}(Z_{k}^{\prime}=0)\leq e^{-A/2^{13}}$ in the right-hand side of (24). To bound the other term in (24) we apply Lemma 4.3 with $i=1$, $j=k$, so that
$\mathbb{P}\big{(}\tau_{1}^{U}-\tau_{1}^{L}<k\mid\mathcal{F}_{\tau^{L}_{1}}\big{)}\leq\exp\left(-\frac{\hat{\varepsilon}n}{2^{10}}\cdot\frac{1}{k}\right)\leq\exp\left(-\frac{\hat{\varepsilon}n}{2^{10}}\cdot\frac{A}{\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)}\right)$
and the proof is completed with (26). ∎
@@
Set $k\coloneqq\lfloor\varepsilon^{-1}k_{0}/A\rfloor$. By applying Lemma 2.9 we infer, that as long as $\tau_{i}^{L}+t\leq\tau_{i}^{U}$, the binomial process $Z_{t}(U_{\tau_{i}^{L}},\varepsilon/2)$ provides a lower coupling for $U_{\tau_{i}^{L}+t}$. Thus, if we set $Z^{\prime}_{t}\coloneqq Z_{t}(\lceil 2^{-10}\varepsilon n\rceil,\varepsilon/2)$, then under the event ${\cal S}_{i}$ the sought probability is at most
$\mathbb{P}\big{(}\tau_{i}^{U}=\infty,Z_{k}(U_{\tau_{i}^{L}},\varepsilon/2)=0\big{)}\leq\mathbb{P}\big{(}Z^{\prime}_{k}=0\big{)}.$
@@
$\rho=\frac{1}{2}(1+\varepsilon/2)(2\rho-\rho^{2})$
and so $\rho=\varepsilon/(1+\varepsilon/2)$. Now, if $A\leq k_{0}$, then we can simply bound
@@
Next we consider the case $k_{0}<A\leq\varepsilon^{-1}k_{0}$. By Lemma 2.11 and Bernoulli’s inequality $(1+x)^{-N}\geq 1-xN$ for $x\geq-1$ and $N\in\mathbb{N}$ and a little algebra we have

### v3

@@ 1 Introduction
In [1] the authors studied this process on several graphs, and established results on the dispersion time and the maximum distance of any particle from the origin at the point in time of dispersion. A particular case considered in [1] is the behaviour when the underlying graph is the complete graph $K_{n}$ with $n$ vertices. In that case we denote the dispersion time by $T_{n,M}$, since, due to symmetry, the choice of the initial vertex does not matter. The most general results come from considering a lazy variant of the dispersion process, which was shown to disperse more particles in a comparable number of steps. More precisely, in this lazy version any unhappy particle moves with probability $q\in(0,1]$ and stays at its current location with probability $1-q$. The main result of [1] regarding the dispersion time of this variant of the process, which we denote by $T_{n,M}(q)$, is that there are constants $c,C>0$ such that, if $M=(1-q/2-\alpha)n$ where $\alpha>0$ is allowed to depend on $n$, then
$T_{n,M}(q)\leq C(q\alpha)^{-1}\ln n~{}\text{ with probability at least }~{}1-C/n,$ (1)
@@
$T_{n,M}(q)\geq e^{cnq^{3}\alpha^{4}}~{}\text{ with probability at least }~{}1-e^{-cnq^{3}\alpha^{4}}.$ (2)
@@
Our main contribution is a thorough analysis of the dispersion process when $q=1$ and, as above, we assume that $M=(1+\varepsilon)n/2$ and $|\varepsilon|=o(1)$. We establish that the process exhibits three qualitatively different behaviours based on the asymptotics of $\varepsilon$, where, informally speaking, $T_{n,M}$ smoothly changes from $|\varepsilon|^{-1}\ln(\varepsilon^{2}n)$ to $n^{1/2}$ and then to $\varepsilon^{-1}e^{\Theta(\varepsilon^{2}n)}$; in particular, $T_{n,M}=\Theta(n^{1/2})$ whenever $M=n/2+O(\sqrt{n})$ within a scaling window of size $O(\sqrt{n})$. We continue with a detailed and formal description of the dispersion process and state our results.
@@
These results imply that there is critical window around $n/2$ of order $n^{1/2}$, where the dispersion time is of order $n^{1/2}$ as well. Moreover, the dispersion time increases smoothly from logarithmic to $n^{1/2}$ and then to exponential when we gradually increase $\varepsilon$. Apart from these estimates we can also use our main theorems to obtain information about, for example, the expectation of $T_{n,M}$. In particular, Theorem 1.1 guarantees that $T_{n,M}$ has an exponential(-ly thin) upper tail and so $T_{n,M}$ is, after appropriate normalization, integrable; we readily obtain
@@ 1.1 Main Proof Ideas
So, as long as $U_{t}$ is much larger than $|\varepsilon|n$, then $U_{t+1}-U_{t}$ is a large negative number in expectation. In other words, when there are ‘many’ unhappy particles at time $t$, the expected number of $U_{t+1}$ decreases by a considerable amount. However, this is no longer true when $U_{t}=O(|\varepsilon|n)$: then the expected decrease of $U_{t}$ might be tiny, or when $\varepsilon>0$, we can even expect that the number of unhappy particles increases. We will use different methods to study the trajectory of $U_{t}$, depending on the range of $\varepsilon$ and whether we are considering the upper or the lower tail.
@@
Let $\mathbb{N}$ denote the set of positive integers and set $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. Given $k\in\mathbb{N}$, we write $[k]\coloneqq\{1,\dots,k\}$. Given functions $f:\mathbb{N}\mapsto\mathbb{R},g:\mathbb{N}\mapsto\mathbb{R}$, we write either $f\ll g$ or $f=o(g)$ if $f(n)/g(n)\rightarrow 0$ as $n\rightarrow\infty$ and $f=O(g)$ if there is a constant $C>0$ such that $f(n)\leq Cg(n)$ for all large enough $n$. We write $f=\Theta(g)$ if $f=O(g)$ and $f=\Omega(g)$, where the latter is the same as $g=O(f)$. Given random variables $X$ and $Y$, we write $X=_{d}Y$ if $X$ and $Y$ have the same distribution, whereas we write $X\leq_{sd}Y$ if $Y$ stochastically dominates $X$, meaning that $\mathbb{P}(X\geq x)\leq\mathbb{P}(Y\geq x)$ for all $x\in\mathbb{R}$. Moreover, given a $p\geq 0$, some events $A,B$ and a $\sigma$-algebra $\cal F$, when we say that $\mathbb{P}(A\mid{\cal F})\geq p$ on $B$, then we mean that this inequality is satisfied only when $B$ occurs, that is, $\mathbb{1}_{B}\cdot\mathbb{P}(A\mid{\cal F})\geq\mathbb{1}_{B}\cdot p$.
@@ 2.2 Bounds on $\mathbb{E}[U_{t+1}\mid U_{t}]$
where the last step can be verified numerically.
Next we turn to the main case $H_{t}\geq 1$. Using $1+x\leq e^{x}$ for $x\in\mathbb{R}$ and $e^{-x}\leq 1-x+x^{2}/2$ for $x\geq 0$ (in this order) we infer
@@
Moreover, since $H_{t}-1=M-U_{t}-1$, $M=(1+\varepsilon)n/2$ and $|\varepsilon|<1/9$, we immediately obtain
@@
$\mathbb{E}\big{[}U_{t+1}\mid U_{t}\big{]}\leq(1+\varepsilon)U_{t}-\frac{U^{2}_{t}}{n}.$
@@ 2.3 Reducing the Number of Unhappy Particles
Here we show a lemma that provides information on the number of steps required to reduce the number of unhappy particles from some number $\ell\leq M$ to some $h<\ell$.
@@
Let $n\in\mathbb{N}$, $t_{0}\in\mathbb{N}_{0}$, $|\varepsilon|<1$, $M=(1+\varepsilon)n/2\in\mathbb{N}$ and $h\geq\max\{2\varepsilon n,1\}>0$. Set $\tau_{h}\coloneqq\inf\{t\geq t_{0}:U_{t}<h\}$. Then, for any $b>0$ and for every $\ell\geq h$
@@ 2.4 Branching Process Approximation
In this subsection we introduce a process that we will use in order to obtain upper and lower couplings for the process of unhappy particles. For $|\kappa|<1$ and some random variable $N\in\mathbb{N}$ define a stochastic process $(Z_{t}(N,\kappa))_{t\in\mathbb{N}_{0}}$ by
$Z_{0}(N,\kappa)=N\quad\text{and}\quad Z_{t}(N,\kappa)=_{d}2\,\text{Bin}\big{(}Z_{t-1}(N,\kappa),(1+\kappa)/2\big{)},~{}t\in\mathbb{N}.$ (8)
We call $(Z_{t}(N,\kappa))_{t\in\mathbb{N}_{0}}$ the binomial process in this paper. We want to stress that the process is well defined even if $\kappa<0$ (as long as $\kappa>-1$). When considering $Z_{t}(N,\kappa)$ we sometimes drop the dependence on $N$ or on $\kappa$, if it is clear from the context or otherwise convenient. In the next lemma we show how to control the number of unhappy particles at some given step in terms of the binomial process.
@@
To see this, at step $t_{0}+t$, remove the particles in ${\cal U}_{t_{0}+t-1}$ from the graph and put them back one after the other (by choosing for each one of them independently and uniformly at random a vertex) to obtain ${\cal U}_{t_{0}+t}$. Note that, during this process, the probability that the $j$-th particle, where $1\leq j\leq U_{t_{0}+t-1}$, is put on one of the, at most $M$, currently occupied vertices is at most $M/n$. Moreover, if this event occurs, then the number of unhappy particles increases by at most 2; it increases by exactly 2 if the $j$-th particle is put on a vertex that is currently occupied by exactly one other particle. This establishes (9). With this at hand we proceed by showing the statement of the lemma by induction over $t$. Note that, for $t=0$, we have that $U_{t_{0}}=Z_{0}$. Next, suppose that the statement is true for $t-1$, where $t\in\mathbb{N}$. Then, for any $z\in\mathbb{N}_{0}$,
@@
If $U_{t_{0}+t-1}\leq\delta n$ by definition of $\tau$, the last expression is at least $(1+\varepsilon)/2-2\delta\eqqcolon q$. Therefore, by induction on $t$, if $(B_{i,t})_{i,t\geq 1}$ denotes a doubly-infinite sequence of iid $\text{Ber}(q)$ random variables,
@@
In order to study the binomial process $(Z_{t}(N,\kappa))_{t\in\mathbb{N}_{0}}$ from (8) we provide an alternative description of it. We can think of any particle $p$ counted in $Z_{0}$ to be the root of an independent Galton-Watson tree $\mathcal{T}_{p}$, where every individual has $2\text{Ber}((1+\kappa)/2)$ offspring, that is, no offspring with probability $(1-\kappa)/2$ and 2 offspring with probability $(1+\kappa)/2$. Obviously, $Z_{t}$ equals in distribution the total number of individuals in all $\mathcal{T}_{p}$’s at distance $t$ from the root. Let $\mathcal{T}(\kappa)$ be a Galton-Watson tree with the aforementioned distribution, and let $\{\mathcal{T}(\kappa)\leq t\}$ be the event that this branching process has at most $t$ levels (in particular, there are no individuals at distance $t$ from the root). Then we immediately obtain the following observation, tailored to our intended application.
@@
we then obtain (10). The lower bound is established in a similar way. Indeed, we obtain from $y_{k}=\frac{y_{k-1}}{1+\varepsilon}(1-x_{k-1}/2)^{-1}$ (see the first ‘$=$’ in (12)) using $\sum_{j=2}^{\infty}q^{j}\leq 2q^{2}$ for all $q\in[0,1/2]$ that
@@ 3 The Upper Tail
Throughout this section we assume that $M=(1+\varepsilon)n/2$, where $|\varepsilon|=o(1)$, $|\varepsilon|<1/9$ and that, say, $n\geq 100$, so that the statements from the previous section apply. Define
@@
that we will mostly use without further reference. The main goal of this section is to prove the following statement, from which the bounds stated in Theorem 1.1 follow immediately. No attempt was made to optimize the constants.
@@
Set $I_{i}=((i-1)k,ik]$ for all $i\in\mathbb{N}$ so that $\mathbb{N}$ is the union of all those intervals. Note that if $T_{n,M}>Aq^{-1}k$, then necessarily the process of unhappy particles cannot terminate in any of the intervals $I_{i}$ with $i\leq Aq^{-1}$. Denote by $\mathcal{S}_{i}$ the event that the process of unhappy particles does not terminate until the end of $I_{i}$; that is,
@@
So,
$\mathbb{P}(T_{n,M}>Aq^{-1}k)\leq\prod_{1\leq i\leq\lfloor Aq^{-1}\rfloor}\mathbb{P}\big{(}\mathcal{S}_{i}\mid\cap_{r=1}^{i-1}\mathcal{S}_{r}\big{)}.$ (15)
@@
If $k\geq\hat{\varepsilon}^{-1}$ and $q\in\mathbb{R}$ satisfy for any $t_{0}\in\mathbb{N}_{0}$
@@
Let $\tau\geq t_{0}$ be the first time such that $U_{\tau}<16\hat{\varepsilon}n$. By conditioning on $\tau$ and $U_{\tau}$ and using the Markov property of the process once more we obtain from the two observations that
@@
Recall that $U_{t}=0$ is equivalent to $T_{n,M}\leq t$. Also, recall the definition of the binomial process that we introduced in (8). Throughout this proof we let $k\coloneqq\lfloor\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)\rfloor$. By applying Lemma 2.8 for any $1\leq r<16\hat{\varepsilon}n$, we obtain
@@ 4 The Lower Tail
Throughout this section we assume that $M=(1+\varepsilon)n/2$, where $|\varepsilon|=o(1)$, $|\varepsilon|<1/9$ and that, say, $n\geq 100$, so that the statements from the Section 2 apply. Define, like previously,
$\hat{\varepsilon}\coloneqq\max\{|\varepsilon|,en^{-1/2}\}.$
We show the following statement, from which Theorem 1.2 follows immediately.
@@
$\mathbb{P}\big{(}T_{n,M}\leq\hat{\varepsilon}^{-1}k_{0}/(2A)\big{)}\leq\left\{\begin{array}[]{ll}3\exp\left(-\frac{A\hat{\varepsilon}^{2}n}{2^{18}k_{0}}\right)&\mbox{, if }A>k_{0}\\ 5A^{-1}&\mbox{, if }1\leq A\leq k_{0}\end{array}\right.\enspace.$
@@
The next lemma asserts that $\mathcal{A}(b)$ occurs for $b=\exp(O(\hat{\varepsilon}n))$ with (very) high probability, meaning that it is very unlikely for $U_{t+1}$ to drop below $2^{-5}U_{t}$ if $U_{t}>2^{-5}\hat{\varepsilon}n$, for all $t\leq b$.
@@
The second issue arising from using the coupling in Lemma 2.9 between $U_{t}$ and the binomial process is that it only works over multiple time steps when the number of unhappy particles remains small throughout. In particular, the coupling provides good estimates only when the number of unhappy particles remains below $\delta n$, where, say, $\delta=\hat{\varepsilon}/8$. In order to ensure that there are sufficiently many time steps until the number of unhappy particles goes above $\delta n$, we introduce a second barrier at $\hat{\varepsilon}n/32$, where we initiate the coupling. More specifically, we define $\tau^{U}_{0}\coloneqq 0$ and iteratively, for $i\in\mathbb{N}$,
@@
where the $D_{\ell}$’s are iid such that $D_{\ell}=1$ with probability $(1+\varepsilon)/2$ and $D_{\ell}=-1$ with probability $(1-\varepsilon)/2$. Define
$S_{k}:=U_{\tau^{L}_{i}}+\cdots+U_{\tau^{L}_{i}+k-1},\quad k\in\mathbb{N}.$
Since the probability that an unhappy particle remains unhappy after moving to a randomly chosen vertex is at most $M/n=(1+\varepsilon)/2$, we note that $U_{\tau^{L}_{i}+1}\leq_{sd}R_{S_{1}}$ and in general
$U_{\tau^{L}_{i}+k}\leq_{sd}R_{S_{k}}.$
@@
throughout. We assume that $k\geq 1$ and $A\geq 4$, as otherwise there is nothing to show. Note that if $T_{n,M}\leq\tau_{1}^{L}+k$, then either $k\leq\tau_{1}^{U}-\tau_{1}^{L}$ and the process stops within $k$ steps after $\tau_{1}^{L}$, or $k>\tau_{1}^{U}-\tau_{1}^{L}$. Since, by Lemma 2.9, the binomial process $Z_{t}(U_{\tau_{1}^{L}},\varepsilon-\hat{\varepsilon}/2)$ provides a lower coupling for $U_{\tau_{1}^{L}+t}$ until time $\tau_{1}^{U}$, we obtain on ${\cal S}_{1}$
@@
This means that $\mathbb{P}(Z_{k}^{\prime}=0)\leq e^{-A/2^{13}}$ in the right-hand side of (24). To bound the other term in (24) we apply Lemma 4.3 with $i=1$, $j=k$, so that on ${\cal S}_{1}$
$\mathbb{P}\big{(}\tau_{1}^{U}-\tau_{1}^{L}<k\mid\mathcal{F}_{\tau^{L}_{1}}\big{)}\leq\exp\left(-\frac{\hat{\varepsilon}n}{2^{10}}\cdot\frac{1}{k}\right)\leq\exp\left(-\frac{\hat{\varepsilon}n}{2^{10}}\cdot\frac{A}{\hat{\varepsilon}^{-1}\ln(\hat{\varepsilon}^{2}n)}\right).$
Note that since $x\mapsto x/\ln x$ is increasing for $x\geq e^{2}$ we have $\hat{\varepsilon}^{2}n/\ln(\hat{\varepsilon}^{2}n)\geq e^{2}/2,$ completing the proof. ∎
@@
Set $k\coloneqq\lfloor\varepsilon^{-1}k_{0}/A\rfloor$. By applying Lemma 2.9 we infer, that as long as $\tau_{i}^{L}+t\leq\tau_{i}^{U}$, the binomial process $Z_{t}(U_{\tau_{i}^{L}},\varepsilon/2)$ provides a lower coupling for $U_{\tau_{i}^{L}+t}$. Thus, if we set $Z^{\prime}_{t}\coloneqq Z_{t}(\lceil 2^{-10}\varepsilon n\rceil,\varepsilon/2)$, then on the event ${\cal S}_{i}$ the sought probability is at most
$\mathbb{P}\big{(}\tau_{i}^{U}=\infty,Z_{k}(U_{\tau_{i}^{L}},\varepsilon/2)=0\mid\mathcal{F}_{\tau_{i}^{L}}\big{)}\leq\mathbb{P}\big{(}Z^{\prime}_{k}=0\big{)}.$
@@
$\rho=\frac{1}{2}(1+\varepsilon/2)(2\rho-\rho^{2}),$
and so $\rho=\varepsilon/(1+\varepsilon/2)$. Now, if $A\leq k_{0}$, then
@@
Next we consider the case $k_{0}<A\leq\varepsilon^{-1}k_{0}$. By Lemma 2.11 and Bernoulli’s inequality $(1+x)^{-N}\geq 1-xN$ for $x>-1$ and $N\in\mathbb{N}$ and a little algebra we have